HMI powered on the Image Stabilization System and both cameras. The operational heaters set points were increased from 10C to 15C and are controlling the instrument temperature nicely. The instrument is responding nominally.
AIA went from a minimal power state to being fully powered. This primarily meant powering up the Camera Electronic Boxes and the ISSs. With the ISSs powered up one is now able to see Guide Telescope signals, which indicate that the GTs are quite healthy and relatively aligned with one another. The thermal control system is appropriately controlling the AIA thermal environment. Over the next few days the temperatures will be increased to aid in outgassing.
EVE completed their scheduled commissioning procedures without errors. These procedures verified EVE autonomy rule functionality, EVE diagnostic channel packet receipt, all 4 filter wheels are functioning nominally and EVE instrument notification commands.